OpenStudy (anonymous):

Is the system consistent or inconsistent? 2x + 3y = 5 and 4x + 6y = 19

OpenStudy (anonymous):

\[3y=-2x+5, y=-\frac{2}{3}x+\frac{5}{3}\] \[6y=-4x+19, y=-\frac{2}{3}x+\frac{19}{6}\] The lines are parallel, it is inconsistent.